OpenStudy (anonymous):

Please double check my answer... What is the inverse of the given function? y = 7x2 - 3.

OpenStudy (anonymous):

I think the correct answer is B??

OpenStudy (anonymous):

y = 7x^2-3 inv x=7y^2-3 x+3 = 7y^2 (x+3)/7 = y^2 sqrt((x+3)/7) = y so. (+/-) sqrt((x+3)/7) so ..B